OpenStudy (anonymous):

What is the prime factorization's of 150, 225, 333, and 96? Use exponents where possible.

OpenStudy (anonymous):

150 = 15 x 10 = 3 x 5 x 2 x 5 = 2 x 3 x 5^2

OpenStudy (klimenkov):

\(150=15\cdot10=3\cdot5\cdot2\cdot5=2\cdot3\cdot5^2\) Watch the analogy.

OpenStudy (anonymous):

225 = 5 x 5 x 3 x 3 = 3^2 x 5^2

OpenStudy (anonymous):

Ok... Thanks :)

OpenStudy (anonymous):

200 = 2 x 100 = 2 x 10 x 10 = 2 x 2 x 5 x 2 x 5 = 2^3 x 5^2

OpenStudy (anonymous):

sorry 200 is actually 333 i was looking at the wrong one

OpenStudy (anonymous):

333 = 3 x 111 = 3 x 3 x 37 = 3^2 x 37

OpenStudy (anonymous):

96 = 2 x 48 = 2 x 2 x 2 x 2 x 3 = 2^4 x 3